A biker's speed is 1.5 time faster than a walker's speed. They start moving simultaneously in the same direction from the same point, and after 1.5 hours, the distance between them is 12.5 miles. What are their speeds?
Answer by lwsshak3(11628)

A biker's speed is 1.5 time faster than a walker's speed. They start moving simultaneously in the same direction from the same point, and after 1.5 hours, the distance between them is 12.5 miles. What are their speeds?
***
let x=walker's speed
1.5x=biker's speed
1.5x-x=.5x=speed by which distance is formed between biker and walker.
speed=distance/travel time
.5x=12.5/1.5=
x=16 2/3
1.5x=25
walker's speed=16 2/3 mph
biker's speed=25 mph